### v2

Abstract
A graph $H$ is an induced minor of a graph $G$ if $H$ can be obtained from $G$ by a sequence of edge contractions and vertex deletions. Otherwise, $G$ is $H$-induced minor-free. In this paper, we provide a different proof of the fact that $K_{2,3}$-induced minor-free graphs admit a quasi-isometry with additive distortion to graphs with tree-width at most two. Our proof yields a $O(nm)$-time algorithm which takes as input a $K_{2,3}$-induced minor-free graph with $n$ vertices and $m$ edges, and outputs a tree-width two graph $H$ with the desired additive distortion. For universally signable graphs, a subclass of $K_{2,3}$-induced minor-free graphs, the time complexity of our algorithm is linear. As a consequence, we obtain a truly sub-quadratic time additive constant factor approximation algorithm to compute the diameter of a universally signable graph. In contrast, assuming the Strong Exponential Time Hypothesis (SETH), the diameter of split graphs (a very restricted class of universally signable graphs), cannot be computed in truly sub-quadratic time [Borassi et al. (ENTCS, 2016)].
1 Introduction
When is a graph quasi-isometric (Definition˜1) to a much simpler host graph? This question is central to coarse graph theory [22]. When the host graphs are restricted to the class of graphs with bounded tree-width, remarkably, the answer is provided by Nguyen, Scott & Seymour [26] and independently by Hickingbotham [24]. Specifically, they proved the following:
Theorem 1 ([26, 24]).
A graph is quasi-isometric to a graph with bounded tree-width if and only if it has a tree-decomposition where each bag consists of a bounded number of balls of bounded diameter.
When quasi-isometric maps with only additive distortions (Definition˜2) are allowed, the situation is much less understood. Brandstädt et al. [6] proved that chordal graphs (i.e., graphs without any induced cycle of length greater than three) admit a quasi-isometry to trees with additive distortion at most $2$. This result has been later generalized to $k$-chordal graphs [12], and more generally to graphs of bounded treelength [18]. Berger & Seymour [3] characterized graphs admitting quasi-isometry with additive distortion to trees. See [19] for a comprehensive overview of graphs admitting a quasi-isometry with additive distortion to trees.
When does a graph class admit quasi-isometry with additive distortion to graphs with tree-width at most two? The above question also relates to the following conjecture of Nguyen, Scott & Seymour [26].
Conjecture 2 ([26]).
There is a constant $k$, such that if a graph $G$ admits a quasi-isometry to a graph of tree-width at most two, then $G$ admits a quasi-isometry with additive distortion to a graph with tree-width at most $k$.
In this paper, we consider the class of $K_{2,3}$-induced minor-free graphs. A graph $H$ is an induced minor of a graph $G$ if $H$ can be obtained from $G$ by a sequence of edge contractions and vertex deletions. Otherwise, $G$ is $H$-induced minor-free. The class of $K_{2,3}$-induced minor-free graphs have been studied extensively in structural graph theory [16, 17]. $K_{2,3}$-induced minor-free graphs generalize many popular graph classes. Examples include outerplanar graphs, circular-arc graphs [20], chordal graphs, universally signable graph [15] etc. Coarse generalizations of $K_{2,3}$-induced minor-free graphs have been studied [13]. A recent result of Albrechtsen et al. [2] implies that $K_{2,3}$-induced minor-free graphs are quasi-isometric to graphs with tree-width at most two. Combining results of Fujiwara & Papasoglu [21] and Nguyen, Seymour & Scott [26] prove that ˜2 holds for $K_{2,3}$-induced minor-free graphs [^2] We thank Robert Hickingbotham for informing us of this connection. In this paper, we provide a different constructive proof for the above. Specifically, we prove the following theorem.
Theorem 3.
There is a constant $c$, and an $O(nm)$-time algorithm that takes as input a $K_{2,3}$-induced minor-free graph $G$ with $n$ vertices and $m$ edges, and outputs a graph $H$ with $V(G)=V(H)$, tree-width at most two, and $\left|d_{G}\left(u,v\right)-d_{H}\left(u,v\right)\right|\leq c$ for any $u,v\in V(G)$.
In the above theorem, $d_{G}\left(.,.\right)$ and $d_{H}\left(.,.\right)$ denote the standard graph metric in $G$ and $H$, respectively. When the input graphs are restricted to universally signable graphs [15], a subclass of $K_{2,3}$-induced minor-free graph, our algorithm runs in linear time, which has the following consequence.
The diameter of a graph is the length of a longest isometric path (i.e. a path of minimum length between the end-vertices) between two of its vertices. Given an undirected graph $G$, Diameter is the algorithmic task of computing the diameter of $G$. Diameter is a fundamental problem with countless applications in computer science. The fastest combinatorial algorithm for Diameter has a running time of $O(nm)$ in $m$-edge, $n$-vertex graphs. Surprisingly, Roditty and Vassilevska Williams [27] proved that a truly sub-quadratic $\left(\frac{3}{2}-\epsilon\right)$-approximation algorithm for Diameter would falsify the Strong Exponential Time Hypothesis [25]. This motivated researchers to study the approximability of Diameter on special graph classes. However, Borassi, Crescenzi & Habib [4] proved that assuming SETH, diameter cannot be computed optimally in truly sub-quadratic time even on split graphs, i.e. graphs whose vertex set can be partitioned into a clique and an independent set. Split graphs are universally signable graphs. Hence assuming SETH, the diameter of universally signable graphs cannot be computed optimally in truly sub-quadratic time. The diameter of graphs with tree-width at most two can be computed in truly sub-quadratic time [7]. Hence, Theorem˜3 imply the following corollary.
Corollary 1.
Diameter admits a truly sub-quadratic additive constant factor approximation algorithm on universally signable graphs.
Below, we provide an informal description of the tools used in the proofs of Theorem˜3. Then, we provide an overview of the proofs.
Important tools: The proof of Theorem˜3 uses layering partition [6, 12] and the notion of strong isometric path complexity [8] of graphs. Layering partitions have been used to solve numerous problems in metric graph theory [6, 12, 13, 19]. See Section˜2.2 for definitions. Informally, in a layering partition, the vertex set of a graph is partitioned into clusters w.r.t a fixed root vertex. Two vertices are in the same cluster if they are at a distance $k$ from the root for some $k\geq 0$ and are end-vertices of a path that is disjoint from the ball of radius k1k-1 centered at the root vertex. It was shown in [6, 12] that the adjacency relation between the clusters is captured by a tree, known as the layering tree. Moreover, for a fixed root vertex, the corresponding layering tree can be constructed in linear time [12].
The strong isometric path complexity is a recently introduced graph invariant that captures how arbitrary isometric paths of a graph can be viewed as a union of few “rooted" isometric paths (i.e. isometric paths with a common end-vertex). See Definition˜4. Informally, if the strong isometric path complexity of a graph $G$ is at most $k$, then for any isometric path $P$ of $G$ and any vertex of $v\in V(G)$, the vertices of $P$ can be covered with $k$-many $v$-rooted isometric paths of $G$. Originally, strong isometric path complexity and its weaker variant isometric path complexity were introduced to propose approximation algorithms for the problem Isometric Path Cover, where the objective is to find a minimum-cardinality set of isometric paths so that each vertex of the graph belongs to at least one of the solution paths. Isometric Path Cover has been studied recently in the algorithmic graph theory community [9, 8, 11]. It was shown that Isometric Path Cover admit a polynomial-time $c$-approximation algorithm on graphs with strong isometric path complexity at most $c$ [8] (while being NP-hard even on chordal graphs whose strong isometric path complexity is at most four [9, 8]). From the structural perspective, many important graph classes such as hyperbolic graphs [23], outerstring graphs [28, 5], and (theta,prism,pyramid)-free graphs [1, 14] have bounded strong isometric path complexity [8]. It was shown in [8, 10] that $K_{2,3}$-induced minor-free graphs have bounded strong isometric path complexity. We shall use this fact in the proof of Theorem˜3. On the other hand, graphs with tree-width at most two have unbounded strong isometric path complexity.
Overview of the proofs of Theorem˜3: Given a $K_{2,3}$-induced minor free graph $G$, we choose a vertex $r\in V(G)$ arbitrarily as the root and construct the layering tree w.r.t $r$, denoted by $\Gamma\left({r},{G}\right)$. Let $R=\{r\}$ be the root cluster. When $\Gamma\left({r},{G}\right)$ is rooted at $R$, there is a natural parent-child relationship among the clusters. For a cluster $S$, the parent set of $S$, denoted as $P\left(r,S\right)$, is the set of vertices that lie in the parent of $S$ and are adjacent to the vertices in $S$. We observe that given a cluster $S$ whose parent is not $R$, $P\left(r,S\right)$ is a minimal cutset of the original input graph. The size of the largest stable set in the graph induced by a minimal cutset of $K_{2,3}$-induced minor-free graph is at most two [17]. Hence, for any cluster $S$ whose parent is not $R$, the size of the largest stable set in the graph induced by $P\left(r,S\right)$ in $G$ is at most two.
We begin our construction by initializing a graph $H=(V(G),\emptyset)$. We traverse $\Gamma\left({r},{G}\right)$ (rooted at $R$) bottom-up and process the clusters in $G$ accordingly. For each cluster $S$, depending on the structure of the graph induced by $P\left(r,S\right)$, we introduce edges in $H$ between the vertices of $S$ and at most two special vertices of $P\left(r,S\right)$. See Algorithm˜1 for details of the construction of $H$. Note that the final resulting graph $H$ may not be a subgraph of $G$.
To prove the upper bound on the distortion, we show that $H$ is connected, and for any edge $uv\in E(G)$, $d_{H}\left(u,v\right)\leq 16$. Next, we show that for any $(u,v)$-isometric path, which is a subpath of some $r$-rooted isometric path in $G$, we have $d_{H}\left(u,v\right)\leq d_{G}\left(u,v\right)+20$. Now consider any $(x,y)$-isometric path $P$ in $G$. Since the strong isometric path complexity of $G$ is bounded, the edges of $P$ can be partitioned into a bounded number of edges and a bounded number of isometric paths that are subpaths of $r$-rooted isometric paths in $G$. Now, using our earlier observations, we show that there exists a constant $c$ (independent of $G$), such that $d_{H}\left(x,y\right)\leq d_{G}\left(x,y\right)+c$. Using a similar technique, we show that for any $u,v\in V(G)$, $d_{G}\left(u,v\right)\leq d_{H}\left(u,v\right)+c$.
To prove the upper bound on the tree-width of $H$, we show that $H$ is triangle-free and does not contain wheel or theta as induced subgraphs. See Section˜2.1 for definitions. The absence of these structures in $H$ implies that $H$ is a universally signable graph, and has tree-width at most two. This concludes the proof of Theorem˜3.
Organisation: In Section˜2, we introduce some notations, recall some existing results, and prove some preliminary observations. In Section˜3, we prove Theorem˜3 and conclude in Section˜4.
2 Preliminaries and some observations
All graphs in this paper are finite, undirected, unweighted, and simple. Let $G$ be a graph. For a set $S\subseteq V(G)$, $G\left[S\right]$ denotes the subgraph of $G$ induced by $S$, and $G-S=G\left[V(G)\setminus S\right]$. When $S=\{w\}$, we write $G-\{w\}$ as GwG-w. For two graphs $G,H$, we write $H\subseteq G$, if $H$ is a subgraph of $G$. A stable set of a graph is a set of pairwise non-adjacent vertices. A cutset of a graph $G$ is a set of vertices of $G$ whose removal disconnects $G$. For two vertices $u,v\in V(G)$, a $(u,v)$-cutset of a graph $G$ is a set of vertices of $G$ whose removal disconnects $u,v$ in $G$. A cutset $S$ of a graph $G$ is minimal if no proper subset of $S$ is a cutset of $G$. Similarly, for two vertices $u,v\in V(G)$, a $(u,v)$-cutset $S$ of a graph $G$ is minimal if no proper subset of $S$ is a $(u,v)$-cutset of $G$.
We state the definition of tree-decompositions for completeness. A tree-decomposition of a graph $G$ is a rooted tree $T$ where each node $v$ is associated to a subset $X_{v}$ of $V(G)$ called bag, such that the set of nodes of $T$ containing a given vertex of $G$ forms a nonempty connected subtree of $T$, and any two adjacent vertices of $G$ appear in a common node of $T$. The width of $T$ is the maximum cardinality of a bag minus $1$. The tree-width of $G$ is the minimum integer $k$ such that $G$ has a tree-decomposition of width $k$.
2.1 $K_{2,3}$-induced minor-free graphs
We shall use the following result of Dallard, Milanič, and Štorgel [17].
Proposition 1 ([17]).
Let $G$ be a $K_{2,3}$-induced minor-free graph. For any two vertices $u,v\in V(G)$, the size of the largest stable set of the graph induced by any minimal $(u,v)$-cutset of $G$ is at most two.
$K_{2,3}$-induced minor-free graphs can be also characterized by forbidding the Truemper configurations and its variants. A hole is a cycle of length at least four. For an integer $t\geq 1$, a prism is a graph made of three vertex-disjoint induced paths $P_{1}\coloneq a_{1}\ldots b_{1}$, $P_{2}\coloneq a_{2}\ldots b_{2}$, $P_{3}\coloneq a_{3}\ldots b_{3}$ of lengths, such that $a_{1}a_{2}a_{3}$ and $b_{1}b_{2}b_{3}$ are triangles and no edges exist between the paths except those of the two triangles. A prism is long if at least one of its three paths has a length of at least 2. A pyramid is a graph made of three induced paths $P_{1}\coloneq a\ldots b_{1}$, $P_{2}\coloneq a\ldots b_{2}$, $P_{3}\coloneq a\ldots b_{3}$, such that $(i)$ they are pairwise vertex-disjoint except at $a$, $(ii)$ two of which are of lengths at least two, $(iii)$ $b_{1}b_{2}b_{3}$ is a triangle, and $(iv)$ no edges exist between the paths except those of the triangle and the three edges incident to $a$. A theta is a graph made of three internally vertex-disjoint induced paths $P_{1}\coloneq a\ldots b$, $P_{2}\coloneq a\ldots b$, $P_{3}\coloneq a\ldots b$ of lengths at least two, and such that no edges exist between the paths except the three edges incident to $a$ and the three edges incident to $b$. A wheel $(H,x)$ consists of a hole $H$, called the rim, and a vertex $x$, called the center, that has at least three neighbors on $H$. A sector of a wheel $(H,x)$ is a path $P$ that is contained in $H$, whose end-vertices are neighbors of $x$ and whose internal vertices are not. Every wheel has at least three sectors. A wheel is broken if at least two of its sectors have a length of at least 2. Dallard et al. [16] proved the following theorem.
Proposition 2 ([16]).
A graph is $K_{2,3}$-induced minor-free if and only if it does not contain a long prism, pyramid, theta, or a broken wheel.
Universally signable graphs form a subclass of $K_{2,3}$-induced minor-free graphs. See [15] for the original definition of universally signable graphs. A graph is universally signable if and only if it does not contain any theta, pyramid, prism, or wheel [15]. Important graph classes like chordal graphs, outerplanar graphs are sub classes of universally signable graphs. Any minimal cutset in a universally signable graph G is either a clique or a stable set of size two [15]. The next proposition follows from the proof of the above result.
Proposition 3 ([15]).
Let $C$ be a cutset of a universally signable graph $G$ that disconnects $G$ into two connected components $G_{1}$ and $G_{2}$ and for each $i\in\{1,2\}$ and each $v\in C$, there is a vertex $u_{i}\in V(G_{i})$ with $vu_{i}\in E(G)$. Then $G\left[C\right]$ is either a clique or a stable set of size two. Moreover, if $G\left[C\right]$ is a stable set of size two, then $V(G)=V(G_{1})\cup V(G_{2})\cup C$.
Proposition˜3 implies the following.
Proposition 4.
Let $G$ be a universally signable graph and $C$ be an induced cycle of $G$ with length greater than three. Let $u,v\in C$ be two non-adjacent vertices of $P$. Then for any indcued path $P$ between $u$ and $v$, we have $V(P)\subseteq V(C)$.
Proposition 5.
The tree-width of a triangle-free universally signable graph is at most two.
2.2 Paths, distances and layering partition
Let $G$ be a graph. The length of a path $P$ is the number of edges in $P$. For two vertices $u,v\in V(G)$, a $(u,v)$-path (resp. $(u,v)$-induced path) is a path (resp. an induced path) between $u$ and $v$. An $(u,v)$-isometric path is an $(u,v)$-path with smallest possible length. The distance between two vertices $u,v\in V(G)$, denoted by $d_{G}\left(u,v\right)$, is the length of an $(u,v)$-isometric path in $G$. For a set $S$ and a vertex $u$, the distance between $u$ and $S$, denoted by $d_{G}\left(u,S\right)=\min\{d_{G}\left(u,v\right)\colon v\in S\}$. When $G$ is clear from the context, we use the notations $d\left(u,v\right)$ and $d\left(u,S\right)$.
Definition 1 ([22]).
An $(M,A)$-quasi-isometry between graphs $G$ and $H$ is a map $f\colon V(G)\rightarrow V(H)$ such that the following holds for fixed constants $M\geq 1,A\geq 0$, (a) $M^{-1}d\left(x,y\right)-A\leq d\left(f(x),f(y)\right)\leq Md\left(x,y\right)+A$for every $x,y\in V(G)$; and (b) for every $z\in V(H)$ there is $x\in V(G)$ such that $d\left(z,f(x)\right)\leq A$. The graphs $G$ and $H$ are quasi-isometric if there exists a quasi-isometric map $f$ between $V(G)$ and $V(H)$.
We highlight the definition of quasi-isometry with additive distortion.
Definition 2.
A class $\mathcal{G}$ of graphs admit a quasi-isometry with additive distortion to a class $\mathcal{H}$ if there exists a constant $A$ (depending only on $\mathcal{G}$ and $\mathcal{H}$) such that any $G\in\mathcal{G}$ admit a $(1,A)$-quasi-isometry to a graph in $\mathcal{H}$.
For an integer $k\geq 0,r\in V(G)$, let $B_{k}\left({r}\right)$ denote the ball of radius $k$ centered at $r$, i.e., $B_{k}\left({r}\right)\coloneq\{u\in G\colon d\left(r,u\right)\leq k\}$. Similarly let $N_{k}(r)$ denote the $k^{th}$ neighborhood of $r$, i.e., $N_{k}(r)\coloneq\{u\in G\colon d\left(r,u\right)=k\}$. For a set of vertices $S\subseteq V(G)$, $B_{k}\left({S}\right)\coloneq\displaystyle\bigcup\limits_{v\in S}B_{k}\left({v}\right)$, and $N_{k}(S)\coloneq\displaystyle\bigcup\limits_{v\in S}N_{k}(v)$. When $k=1$, we shall write $N_{1}(r)$ and $N_{1}(S)$ simply as $N(r)$ and $N(S)$.
Let $G$ be a graph and $r$ be vertex of $G$. For $k\geq 2$ and two vertices $u,v\in N_{k}(r)$, an $(u,v)$-induced path $P$ is an $(u,v)$-upper path w.r.t $r$ if $V\left(P-\{u,v\}\right)\subseteq B_{k-1}\left({r}\right)$. Since $B_{k-1}\left({r}\right)\cup\{u,v\}$ induces a connected subgraph of $G$, a $(u,v)$-upper path w.r.t $r$ always exists.
We recall the notion of layering partition, introduced in [6, 12]. Let $G$ be a graph and $r$ be any vertex of $G$. A subset $S\subseteq V(G)$ is a cluster w.r.t $r$ if $S$ is the maximal subset such that (i) there exists an integer $k\geq 1$ with $S\subseteq N_{k}(r)$; (ii) any two vertices $x,y\in S$ lie in the same connected component in $G-B_{k-1}\left({r}\right)$. Let $G$ be a graph, $r$ be a vertex of $G$ and $S$ be a cluster such that $S\subseteq N_{k}(r)$ for some integer $k\geq 1$. The parent set of $S$, denoted by $P\left(r,S\right)$, is the set $\left\{v\in N_{k-1}(r)\colon uv\in E(G),u\in S\right\}$. Observe that $P\left(r,S\right)$ is a subset of some cluster w.r.t $r$.
Let $G$ be a graph and $r$ be a vertex of $G$. Two clusters $C,C^{\prime}$ w.r.t $r$ are adjacent if there is an edge in $G$ between some $x\in C$ and $y\in C^{\prime}$. For a graph $G$ and a vertex $r$, let $\Gamma\left({r},{G}\right)$ be the graph whose vertices are the maximal clusters w.r.t $r$, and two vertices are adjacent if the corresponding clusters are adjacent.
Proposition 6 ([12, 6]).
For any graph $G$ and a vertex $r\in V(G)$, $\Gamma\left({r},{G}\right)$ is a tree.
Below, we show that the parent set of every cluster is a minimal $(r,v)$-cutset where $v$ is a vertex of the cluster.
Lemma 4.
Let $G$ be a graph, $r$ be a vertex, $S$ be a cluster w.r.t $r$ and $u\in S$ such that $P\left(r,S\right)\neq\{r\}$. Then $P\left(r,S\right)$ is a minimal $(r,u)$-cutset of $G$.
Proof.
Let $X\coloneq P\left(r,S\right)$ and $S\subseteq N_{k}(r)$ with $k\geq 2$. For any $u\in S$, an $(r,s)$-path in $G$ must contain a vertex of $X$, and therefore, $r$ and $u$ lie in different connected components in GXG-X. Hence, $X$ is a cutset.
Let $Y\subset X$, H=GYH=G-Y, $u\in X\setminus Y$, and $v\in S$ with $uv\in E(H)$. Let $Z_{1}$ be the set of vertices that lie in $S$ or in some cluster which is a descendent of $S$ in $\Gamma\left({r},{G}\right)$. Observe that $Z_{1}\subset V(H)$ and in particular, $v\in Z_{1}$. Since any vertex $w\in Z_{1}$ is connected to $v$ by a path in $H\left[Z_{1}\right]$, we have that $H\left[Z_{1}\right]$ is connected. Let $Z_{2}=V(G-Z_{1}-Y)$. By definition, $V(H)=Z_{1}\cup Z_{2},r\in Z_{2}$ and $u\in Z_{2}$. Observe that for any vertex $w\in Z_{2}$, there is a $(r,w)$-path in $H$. Hence $H\left[Z_{2}\right]$ is connected. As $uv\in E(H)$, $H$ is connected. ∎
Lemma˜4 and Proposition˜1 imply the following properties of clusters of $K_{2,3}$-induced minor-free graphs.
Lemma 5.
Let $G$ be a $K_{2,3}$-induced minor-free graphs, $r$ be a vertex, $S$ be a cluster w.r.t $r$, and $F\coloneq G\left[P\left(r,S\right)\right]$. Then all of the following holds. (a) FFhas at most two connected components, (b) If $F$ is connected then it has diameter at most three, (c) If $F$ has two connected components, then each connected component is a clique.
For technical reasons, we will need the following lemma.
Lemma 6.
Let $G$ be a $K_{2,3}$-induced minor-free graph, $r$ be a vertex, and $S$ be a cluster w.r.t $r$, $S_{1},S_{2}$ be two children of $S$ in $\Gamma\left({r},{G}\right)$ such that for $i\in\{1,2\}$, $G\left[P\left(r,S_{i}\right)\right]$ contains two connected components $C_{i}$ and $D_{i}$. Let $\mathcal{X}=\{(Y,Z)\in\{C_{1},D_{1}\}\times\{C_{2},D_{2}\}\colon V(Y)\cap V(Z)\neq\emptyset\}$. Then $\left|\mathcal{X}\right|\leq 1$.
Proof.
If there exists a connected component $Y\in\{C_{1},D_{1}\}$ containing vertices from both $C_{2}$ and $D_{2}$, then due to Lemma˜5(c), there will be an edge in $G\left[P\left(r,S_{2}\right)\right]$ between vertices of $C_{2}$ and $D_{2}$, which is a contradiction. Now assume $|\mathcal{X}|\geq 2$. Without loss of generality, let $C_{1}$ contains a vertex $c$ from $C_{2}$. Then, $D_{1}$ must contain a vertex $d$ from $D_{2}$. For $i\in\{1,2\}$, let $c_{i}\in N(c)\cap S_{i}$ and $d_{i}\in N(d)\cap S_{i}$. Let $k=d\left(r,S\right)$. For $i\in\{1,2\}$, let $P_{i}$ denote a $(c_{i},d_{i})$-induced path in $G-B_{r}\left({k}\right)$. (Since $c_{i},d_{i}$ lie in the same cluster in $G$, $P_{i}$ exists). Let $P$ be an $(c,d)$-upper path w.r.t $r$. the graph induced by the vertices of $P,P_{1},P_{2}$, contains a theta, which contradicts Proposition˜2. ∎
2.3 Strong isometric path complexity
Given a graph $G$ and a vertex $v$ of $G$, a set $S$ of isometric paths of $G$ is $v$-rooted if $v$ is one of the end-vertices of all the isometric paths in $S$. A path $P$ is covered by a set $\mathcal{Q}$ of paths if $V(P)\subseteq\displaystyle\bigcup\limits_{Q\in\mathcal{Q}}V(Q)$.
Definition 3 ([8, 10]).
The “isometric path complexity” of a connected graph $G$, denoted by $ipco\left(G\right)$, is the minimum integer $k$ such that there exists a vertex $v\in V(G)$ that satisfy the following property: any isometric path of $G$ can be covered by $k$ many $v$-rooted isometric paths.
Definition 4 ([8, 10]).
The “strong isometric path complexity” of a connected graph $G$, denoted by $sipco\left(G\right)$, is the minimum integer $k$ such that all vertices of $v\in V(G)$ satisfy the following property: any isometric path of $G$ can be covered by $k$ many $v$-rooted isometric paths.
A result of [8] implies that the strong isometric path complexity of graphs without a theta, pyramid, or a long prism is at most $71$. Now Proposition˜2, imply the following:
Proposition 7.
The strong isometric path complexity of $K_{2,3}$-induced minor-free graphs is at most $71$.
Below, we show how the notion of strong isometric path complexity can be used to prove quasi-isometry with additive distortion between graphs.
[…]
Figure 1: Illustration of the proof of Lemma˜7. Here $z=3$. The dashed curve from $u$ to $v$ indicates a $(u,v)$-path in $Y$. The solid black piecewise linear curve between $(u,v)$ that doesn’t pass through $r$, indicates the $(u,v)$-isometric path $P$ in $X$. For $p\in\{a_{1}=u,b_{2},b_{3}=v\}$, the strictly monotone solid black curves between $r,p$ indicate a $(r,p)$-isometric path in $X$.
Lemma 7.
Let $X$ and $Y$ be two graphs such that $V(X)=V(Y)$, and $r\in V(X)$ be a vertex. Let $t\geq 1,k\geq 1$ be two integers such that the following holds.
  1. (i)
    for any edge $ab\in E(X)$, $d_{Y}\left(a,b\right)\leq k+1$;
  2. (ii)
    for any $r$-rooted isometric path $Q\subseteq X$ and $\{a,b\}\subseteq V(Q)$, $d_{Y}\left(a,b\right)\leq d_{X}\left(a,b\right)+t$.
Then, for $\{u,v\}\subseteq V(X)$, $d_{Y}\left(u,v\right)\leq d_{X}\left(u,v\right)+t\cdot sipco\left(X\right)+k(sipco\left(X\right)-1)$.
Proof.
Let $P$ be any $(u,v)$-isometric path in $X$ and $\mathcal{Q}$ be a minimum cardinality set of $r$-rooted isometric paths in $X$ that covers $P$ and for any $Q\in\mathcal{Q}$, $|V(P)\cap V(Q)|$ is maximized. See Figure˜1 for illustration. Let $z\coloneq|\mathcal{Q}|$ and by definition, $z\leq sipco\left(X\right)$. Observe that for any $Q\in\mathcal{Q}$, the path induced by $V(P)\cap V(Q)$ is an isometric path in $X$. Moreover, there exists exactly one path $Q_{1}\in\mathcal{Q}$ that contains $u$. Otherwise, if there exists a path $Q^{\prime}\in\mathcal{Q}\setminus\{Q_{1}\}$ containing $u$, then either $V(P)\cap V(Q^{\prime})$ contains $V(P)\cap V(Q_{1})$, or vice versa. But this contradicts the minimality of $\mathcal{Q}$.
Let $Q_{1}\in\mathcal{Q}$ be the path containing $u$, and define $P_{1}\coloneq P\left[V(P)\cap V(Q_{1})\right]$. Let $a_{1}\coloneq u$, and $b_{1}$ be the end-vertex of $P_{1}$ distinct from $a_{1}$. For $i\in\{2,\ldots,z\}$, $S_{i}\coloneq V(P)\setminus\left\{\displaystyle\bigcup\limits_{j<i}V(P_{j})\right\}$, $a_{i}\in S_{i}$ be the vertex adjacent to $b_{i-1}$ in $P$, $Q_{i}\in\mathcal{Q}$ be the path containing $a_{i}$, and $P_{i}\coloneq P\left[S_{i}\cap V(Q_{i})\right]$. Note that $b_{z}=v$. Hence,
$\begin{split}d_{Y}\left(u,v\right)&=d_{Y}\left(a_{1},b_{z}\right)\\ &\leq\displaystyle\sum\limits_{i=1}^{z}d_{Y}\left(a_{i},b_{i}\right)+\displaystyle\sum\limits_{i=1}^{z-1}d_{Y}\left(b_{i},a_{i+1}\right)\\ &\leq\displaystyle\sum\limits_{i=1}^{z}(d_{X}\left(a_{i},b_{i}\right)+t)+\displaystyle\sum\limits_{i=1}^{z-1}(k+1)\hskip 10.0pt\\ &\leq\displaystyle\sum\limits_{i=1}^{z}d_{X}\left(a_{i},b_{i}\right)+\displaystyle\sum\limits_{i=1}^{z-1}d_{X}\left(b_{i},a_{i+1}\right)+t\cdot z+k(z-1)\\ &\leq d_{X}\left(a_{1},b_{z}\right)+t\cdot z+k(z-1)\\ &=d_{X}\left(u,v\right)+t\cdot z+k(z-1)\end{split}$
This completes the proof. ∎
3 Proof of Theorem˜3
For the entirety of this section, $G$ shall denote a fixed $K_{2,3}$-induced minor-free graph, and $r$ shall denote a fixed vertex of $G$. All clusters (and their parent sets) are defined w.r.t the metric in $G$ and w.r.t the vertex $r$. The output of Algorithm˜1 with $G$ and $r$ as input will be denoted as $H$. See Figure˜2 for an example. The sets $N_{k}(r)$, $B_{k}\left({r}\right)$, $P\left(r,S\right)$ always contains vertices of $G$ and are defined w.r.t the metric in $G$. In Section˜3.1, we observe some properties of $H$. In Section˜3.2, we prove lemmas relating the metric of $G$ and $H$. In Section˜3.3, we prove some properties on the induced cycles of $H$, which we use in Section˜3.4 to show that $H$ has tree-width at most two. We complete the proof in Section˜3.5.
$r$
(a) A graph $G$.
$r$
(b) The graph $H$.
Figure 2: The shaded regions indicate the clusters w.r.t $r$. The graph $H$ is the output of Algorithm˜1 with $G$ and $r$ as input.
Input : A $K_{2,3}$-induced minor-free graph $G$ and a vertex $r$.
Output : A graph $H=(V(G),E^{\prime})$.
1
21mm
3
4Let $E^{\prime}=\emptyset$.
5Construct $T=\Gamma\left({r},{G}\right)$.
6for each cluster $S\in V(T)$ do
7 if $G\left[P\left(r,S\right)\right]$ is connected then
8      Choose a vertex $w$ arbitrarily in $P\left(r,S\right)$.
9    for each vertex $v\in S$ do
10       $E^{\prime}=E^{\prime}\cup\{vw\}$.
11
12 else
13      Let $G\left[P\left(r,S\right)\right]$ has two connected components $C_{1},C_{2}$. For $i\in\{1,2\}$, choose a vertex $w_{i}\in C_{i}$ and let $D_{i}\coloneq\{u\in S\colon\exists v\in C_{i},uv\in E(G)\}$.
14    if $D_{1}\cap D_{2}=\emptyset$ then
15       for each vertex $v\in D_{1}$ do
16          $E^{\prime}=E^{\prime}\cup\{vw_{1}\}$.
17       for each vertex $v\in D_{2}$ do
18          $E^{\prime}=E^{\prime}\cup\{vw_{2}\}$.
19       if there exists an edge $uv\in E(G)$ with $u\in D_{1},v\in D_{2}$ then
20          $E^{\prime}=E^{\prime}\cup\{uv\}$
21
22    else
23         Color the vertices of $C_{1}$ as blue.
24       for each vertex $v\in D_{1}$ do
25          $E^{\prime}=E^{\prime}\cup\{vw_{1}\}$.
26       for each vertex $v\in D_{2}\setminus D_{1}$ do
27          $E^{\prime}=E^{\prime}\cup\{vw_{2}\}$.
28         Choose a vertex $u\in D_{1}\cap D_{2}$ arbitrarily and $E^{\prime}=E^{\prime}\cup\{uw_{2}\}$
29
30
31 return $H=(V(G),E^{\prime})$.
32
Algorithm 1 An algorithm to embed $K_{2,3}$-induced minor-free graph on $K_{4}$-minor free graph.
3.1 Basic facts about $H$
Observe that, for every cluster $S$ of $G$ and vertex $v\in S$, we have that $v$ is adjacent to at least one vertex from $P\left(r,S\right)$ in $H$. Hence, the graph $H$ is connected. Below we state some observations which follow immediately from Algorithm˜1. We provide the proofs for completeness.
Observation 1.
Let $u,v\in V(G)$ be two vertices such that $d_{G}\left(r,v\right)=d_{G}\left(r,u\right)+1$ and $uv\in E(H)$. Let $S$ be the cluster in $G$ that contains $v$. Then (a) $u\in P\left(r,S\right)$; and (b) there exists $w\in V(G)$ such that $vw\in E(G)$ and $u,w$ lies in the same connected component in $G\left[P\left(r,S\right)\right]$.
Proof.
$(a)$ follows immediately from Algorithms˜1 and 1 of Algorithm˜1. To prove $(b)$, first consider the case when $F\coloneq G\left[P\left(r,S\right)\right]$ is connected. Clearly, $v$ has a neighbor $w\in P\left(r,S\right)$ such that $u$ and $w$ lie in the same connected component in $F$. Now consider the case when $F$ has two connected components $C_{1},C_{2}$ and for $i\in\{1,2\}$, let $D_{i}\coloneq N(C_{i})\cap S$. Since $C_{1},C_{2}$ are cliques in $G$ (due to Lemma˜5(c)), and since there exists $i\in\{1,2\}$ such that $u\in C_{i}$, Algorithm˜1 ensures that $v\in D_{i}$. This implies $C_{i}$ has a vertex $w$ with $\{vw,uw\}\subseteq E(G)$. ∎
˜1 immediately implies the following.
Observation 2.
Let $P$ be an $(u,v)$-path in $H$ and $S_{u},S_{v}$ be the clusters containing $u,v$, respectively. Let $S$ be the least common ancestor of $S_{u}$ and $S_{v}$ in $\Gamma\left({r},{G}\right)$. Then at least one vertex of $P$ lies in $S$.
Observation 3.
Let $S$ be a cluster and $F\coloneq\{v\in P\left(r,S\right)\colon\exists u\in S,uv\in E(H)\}$. Then, (a) $|F|\leq 2$; and (b) if $|F|=2$ then $G\left[P\left(r,S\right)\right]$ has two connected components in $G$.
Proof.
The proof follows from Algorithms˜1 and 1 of Algorithm˜1. ∎
Observation 4.
Let $S$ be a cluster and let $F\coloneq\{u\in S\colon\exists\{v_{1},v_{2}\}\subseteq P\left(r,S\right),\{uv_{1},uv_{2}\}\subseteq E(H)\}$. Then $|F|\leq 1$.
Proof.
The proof follows from Algorithm˜1 of Algorithm˜1. ∎
Observation 5.
Let $S$ be a cluster and let $F\coloneq\{uv\in E(H)\colon u,v\in S\}$. Then, (a) $|F|\leq 1$; (b) if $F=\{uv\}$, then $G\left[P\left(r,S\right)\right]$ has two connected components, $u,v$ do not have a common neighbor in $H$, and the set $F^{\prime}=\{w\in S\colon\exists\{v_{1},v_{2}\}\subseteq P\left(r,S\right),\{wv_{1},wv_{2}\}\subseteq E(H)\}$ is empty.
Proof.
The proof follows from Algorithm˜1. ∎
We end this section by showing that the distances between $r$ and any vertex in $G$ remains the same in $H$.
Lemma 8.
For any vertex $u\in V(G)$, $d_{H}\left(r,u\right)=d_{G}\left(r,u\right)$.
Proof.
Observe that $ru\in E(G)$ implies $ru\in E(H)$. Assume that for an integer $k\geq 2$, $d_{G}\left(r,u\right)=k-1$ implies $d_{H}\left(r,u\right)=d_{G}\left(r,u\right)$. Let $v$ be a vertex with $d_{G}\left(r,v\right)=k$. Let $S$ be the cluster that contains $v$. Notice that any $(r,v)$-isometric path in $H$ contains a vertex $u\in P\left(r,S\right)$. Since $P\left(r,S\right)\subseteq N_{k-1}(r)$, we have that $d_{G}\left(r,u\right)=d_{H}\left(r,u\right)=k-1$. Hence, $d_{H}\left(r,v\right)=k$. ∎
3.2 Distances in $G$ and $H$
We show that, two adjacent vertices that lie in the same cluster in $G$, remain close to each other in $H$.
Lemma 9.
For any edge $xy\in E(G)$ with $d_{G}\left(r,x\right)=d_{G}\left(r,y\right)$, $d_{H}\left(x,y\right)\leq 5$.
Proof.
Let $k=d_{G}\left(r,x\right)=d_{G}\left(r,y\right)$, and $S\subseteq N_{k}(r)$ be the cluster containing $u,v$. Due to Lemma˜5, the graph $F\coloneq G\left[P\left(r,S\right)\right]$ has at most two connected components. If $F$ is connected, then consider the vertex $w$ which is chosen in Algorithm˜1 of Algorithm˜1. Clearly, $\{xw,yw\}\subseteq E(H)$ (due to Algorithm˜1 of Algorithm˜1), and therefore $u,v$ has a common neighbor in $H$. Hence, $d_{H}\left(x,y\right)\leq 2$. Suppose $F$ has two connected components $C_{1},C_{2}$. For $i\in\{1,2\}$, let $D_{i}\coloneq\{u\in S\colon\exists v\in C_{i},uv\in E(G)\}$.
Suppose $D_{1}\cap D_{2}=\emptyset$. If there exists a $i\in\{1,2\}$, such that both $x,y\in D_{i}$, then there is a vertex $w\in C_{i}$ such that $\{xw,yw\}\subseteq E(H)$. Otherwise, due to Algorithm˜1, there exists an edge $uv\in E(H)$ such that $u,x$ has a common neighbor in $H$ and $v,y$ has a common neighbor in $H$. Hence, $d_{H}\left(x,y\right)\leq 5$.
Now assume $D_{1}\cap D_{2}\neq\emptyset$. Without loss of generality assume that, the vertices of $C_{1}$ are colored as blue in Algorithm˜1. If $\{x,y\}\subseteq D_{1}$ or $\{x,y\}\subseteq D_{2}\setminus D_{1}$, then $x,y$ has a common neighbor in $H$ (due to Algorithms˜1 and 1 of Algorithm˜1), and hence, $d_{H}\left(x,y\right)=2$. Otherwise, without loss of generality assume $x\in D_{1}$ and $y\in D_{2}\setminus D_{1}$. Due to Algorithm˜1 of Algorithm˜1, there exists a vertex $w\in D_{1}$ such that $x,w$ has a common neighbor in $H$ and $y,w$ has a common neighbor in $H$. Hence $d_{H}\left(x,y\right)\leq 4$. ∎
Next, we extend Lemma˜9 and show that adjacent vertices in $G$ remain close to each other in $H$.
Lemma 10.
For any edge $xy\in E(G)$, $d_{H}\left(x,y\right)\leq 16$.
Proof.
Let $xy\in E(G)$, and without loss of generality assume that $d_{G}\left(r,y\right)\leq d_{G}\left(r,x\right)$. Let $k=d_{G}\left(r,x\right)$, and $S\subseteq N_{k}(r)$ be the cluster containing $x$. If $d_{G}\left(r,y\right)=k$, then $y\in S$ and we are done by Lemma˜9. Now consider the case when $d_{G}\left(r,y\right)=k-1$. By definition, $y\in P\left(r,S\right)$. Due to Lemma˜5, $F\coloneq G\left[P\left(r,S\right)\right]$ has at most two connected components. If $F$ is connected, then let $w$ be the vertex chosen in Algorithm˜1 of Algorithm˜1. Note that $xw\in E(H)$. Due to Lemma˜5(b), there is a $(y,w)$-path $P\subseteq F$ of length at most three. Applying Lemma˜9 on the each edge of the path $P$, we have that $d_{H}\left(x,y\right)\leq 16$.
Suppose $F$ has two connected components $C_{1},C_{2}$. Due to Lemma˜5(c), each of $C_{1}$ and $C_{2}$ are cliques in $G$. For $i\in\{1,2\}$, let $D_{i}\coloneq\{u\in S\colon\exists v\in C_{i},uv\in E(G)\}$. Now we have the following cases.
If $D_{1}\cap D_{2}=\emptyset$, then there exists an integer $i\in\{1,2\}$ such that $x\in D_{i}$ and $y\in C_{i}$. In this case, there exists a vertex $w\in C_{i}$ such that $xw\in E(H)$ (due to Algorithms˜1 and 1). Applying Lemma˜9 on the edge $yw\in E(G)$, we have that $d_{H}\left(x,y\right)\leq 6$.
Consider the case, when $D_{1}\cap D_{2}\neq\emptyset$. Without loss of generality assume that the vertices of $C_{1}$ are colored blue in Algorithm˜1. If $y\in C_{1},x\in D_{1}$ or if $y\in C_{2},x\in D_{2}\setminus D_{1}$, then using analogous arguments as before we have that $d_{H}\left(x,y\right)\leq 6$. Now suppose $y\in C_{2}$ and $x\in D_{1}\cap D_{2}$. Then, due to Algorithm˜1 of Algorithm˜1, there exist vertices $w_{1},v,w_{2}$ such that $\{xw_{1},w_{1}v,vw_{2}\}\subseteq E(H)$ and $w_{2}\in C_{2}$. Now applying Lemma˜9 on the edge $yw_{2}\in E(G)$, we have that $d_{H}\left(x,y\right)\leq 8$. ∎
Note that the graph $H$ (i.e. the output of Algorithm˜1) may not be a subgraph of $G$. In the following lemma, we show that the adjacent vertices in $H$ are also close to each other in $G$.
Lemma 11.
For any $xy\in E(H)$, $d_{G}\left(x,y\right)\leq 4$.
Proof.
Let $xy$ be any edge of $H$. Without loss of generality assume $d_{G}\left(r,y\right)\leq d_{G}\left(r,x\right)$. If $d_{G}\left(r,x\right)=d_{G}\left(r,y\right)$, then Algorithm˜1 implies that $xy\in E(G)$. Hence, the lemma is true in this case. Now assume $d_{G}\left(r,y\right)<d_{G}\left(r,x\right)$. Algorithm˜1 ensures that $d_{G}\left(r,y\right)=d_{G}\left(r,x\right)-1$. Let $S$ be the cluster that contains $x$ and let $F\coloneq G\left[P\left(r,S\right)\right]$. Due to ˜1(a), $y\in V(F)$. Due to ˜1(b), there exists a vertex in $z\in V(F)$ such that $xz\in E(G)$ and $y,z$ lie in the same connected component in $F$. If $F$ is connected, then due to Lemma˜5(b) we have that $d_{G}\left(z,y\right)\leq 3$ and, thus $d_{G}\left(x,y\right)\leq 4$. Otherwise, due to Lemma˜5(c), $F$ must have two connected components both of which are cliques. This implies $yz\in E(G)$ and thus $d_{G}\left(x,y\right)\leq 2$. ∎
In the next lemma, we show that if a pair of vertices lie in a $r$-rooted isometric path in $G$, then their distance in $H$ is increased by an additive factor of at most $20$.
Lemma 12.
For any $r$-rooted isometric path $P\subseteq G$ and two vertices $x,y\in V(P)$, we have $d_{H}\left(x,y\right)\leq d_{G}\left(x,y\right)+20$.
Proof.
Let $Q$ be the $(x,y)$-subpath of $P$. Assume that the length of $Q$ is at least two. (Otherwise, we are done by Lemma˜10.) Without loss of generality assume $d_{G}\left(r,y\right)\leq d_{G}\left(r,x\right)$ and $Q\coloneq x_{\ell}x_{\ell+1}\ldots,x_{k}$, with $x_{\ell}\coloneq y$ and $x_{k}\coloneq x$. The above definitions imply $d_{G}\left(x,y\right)=d_{G}\left(x_{k},x_{\ell}\right)=k$-$\ell$. Observe that any cluster contains at most one vertex of $Q$. For $i\in[\ell,k]$, let $S_{i}$ denote the cluster that contains $x_{i}$. Let $j\in[\ell,k]$ be the minimum integer that satisfy one of the following:
  1. (i)
    $j=k$,
  2. (ii)
    $G\left[P\left(r,S_{j+1}\right)\right]$ is connected,
  3. (iii)
    $G\left[P\left(r,S_{j+1}\right)\right]$ has two connected components $C_{1},C_{2}$ such that the sets $D_{1}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{1},uv\in E(G)\}$ and $D_{2}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{2},uv\in E(G)\}$ have a common vertex.
We prove the following claim.
Claim 12.1.
$d_{H}\left(x_{k},x_{j}\right)\leq k-j+15$.
Proof.
If $k=j$, then the claim is trivially true. So we assume $j<k$. Let $x^{\prime}_{k}\coloneq x_{k}$ and for each $i\in[j,k)$, let $x^{\prime}_{i}\in S_{i}$ denote a vertex such that $x^{\prime}_{i}x^{\prime}_{i+1}\in E(H)$. Let $R\coloneq x^{\prime}_{k}x^{\prime}_{k-1}\ldots x^{\prime}_{j}$. If $G\left[P\left(r,S_{j+1}\right)\right]$ is connected, then due to Lemma˜5(b), there is a $(x^{\prime}_{j},x_{j})$-induced path $Q^{\prime}\subseteq F$ of length at most three in $G$. Applying Lemma˜9, on each edge of $Q^{\prime}$ we have that $d_{H}\left(x^{\prime}_{j},x_{j}\right)\leq 15$ and hence $d_{H}\left(x_{k},x_{j}\right)\leq k-j+15$.
Now suppose $G\left[P\left(r,S_{j+1}\right)\right]$ has two connected components $C_{1},C_{2}$ such that the sets $D_{1}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{1},uv\in E(G)\}$ and $D_{2}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{2},uv\in E(G)\}$ have a common vertex. If there exists $i\in\{1,2\}$ with $x_{j},x^{\prime}_{j}\in C_{i}$, then applying Lemma˜9 on the edge $x_{j}x^{\prime}_{j}\in E(G)$, we have $d_{G}\left(x^{\prime}_{j},x_{j}\right)\leq 5$. Since $D_{1}\cap D_{2}\neq\emptyset$, due to Algorithm˜1, there exist $u\in S_{j+1}$ and $v\in S_{j}$ such that $\{x^{\prime}_{j}u,uv\}\subseteq E(H)$ and $vx_{j}\in E(G)$. Since $\{v,x_{j}\}\subseteq S_{j}$, Lemma˜9 implies $d_{H}\left(x^{\prime}_{j},x_{j}\right)\leq 7$. Hence, $d_{H}\left(x_{k},x_{j}\right)\leq k-j+7$. ∎
Let $x^{\prime}_{j}=x_{j}$ and for each $i\in[\ell,j-1]$, let $x^{\prime}_{i}$ be a neighbor of $x^{\prime}_{i+1}$ in $H$. We prove the following claim.
Claim 12.2.
For each $i\in[\ell,j-1]$, $x_{i}x^{\prime}_{i}\in E(G)$.
Proof.
For each $i\in[\ell,j-1]$, let the two connected components of $G\left[P\left(r,S_{i+1}\right)\right]$ are $C^{1}_{i}$ and $C^{2}_{i}$. Due to Proposition˜1, both $C^{1}_{i}$ and $C^{2}_{i}$ are cliques in $G$. For each $t\in\{1,2\}$, let $D^{t}_{i}\coloneq\{u\in S_{i+1}\colon\exists v\in C^{t}_{i},uv\in E(G)\}$. By definition, $D^{1}_{i}\cap D^{2}_{i}=\emptyset$. Now without loss of generality assume that $x_{i+1}\in D^{1}_{i}$. Therefore, any vertex in $S_{i}$ which is adjacent to $x_{i+1}$ in $G$ must lie in $C^{1}_{i}$. Therefore, $x_{i}\in C^{1}_{i}$. Now due to Algorithm˜1, $x^{\prime}_{i}\in C^{1}_{i}$. Since, $C^{1}_{i}$ is a clique, we have that $x_{i}x^{\prime}_{i}\in E(G)$. ∎
Let $R_{1}$ be a $(x_{k},x_{j})$-isometric path in $H$, define $R_{2}\coloneq x^{\prime}_{j}x^{\prime}_{j-1}\ldots x^{\prime}_{\ell}$ and $R_{3}$ be a $(x^{\prime}_{\ell},x_{\ell})$-isometric path in $H$. Due to ˜12.1, length of $R_{1}$ is at most kj+15k-j+15. Length of $R_{2}$ is at most $j$-$\ell$. Due to ˜12.2 and Lemma˜9, length of $R_{3}$ is at most $5$. Hence, the length of the path $R_{1}\cup R_{2}\cup R_{3}$ is at most $k-\ell+20$. Hence $d_{H}\left(x,y\right)\leq d_{G}\left(x,y\right)+20$. ∎
In the next lemma, we show that if a pair of vertices lie in a $r$-rooted isometric path in $H$, then their distance in $G$ is decreased by an additive factor of at most five. The proof is similar to that of Lemma˜12.
Lemma 13.
For any $r$-rooted isometric path $P\subseteq H$, and two vertices $x,y\in V(P)$, $d_{G}\left(u,v\right)-5\leq d_{H}\left(u,v\right)$.
Proof.
Without loss of generality, assume that $d_{H}\left(r,y\right)\leq d_{H}\left(r,x\right)$. If $xy\in E(H)$, then we are done by Lemma˜11. Otherwise, let $Q$ be the $(x,y)$-subpath of $P$. Let $k\coloneq d_{H}\left(r,x\right)$, $\ell\coloneq d_{H}\left(r,y\right)$, $x_{k}\coloneq x$, $x_{\ell}=y$ and $Q\coloneq x_{k}x_{k-1}\ldots x_{\ell}$. Observe that for $i\in[\ell,k]$, $d_{H}\left(r,x_{i}\right)=i$. Due to Lemma˜8, $d_{G}\left(r,x_{i}\right)=i$. Therefore, any cluster contains at most one vertex of $Q$. For $i\in[\ell,k]$, let $S_{i}$ be the cluster that contains $x_{i}$. Due to ˜1(a), for $i\in[\ell+1,k]$, $x_{i-1}\in P\left(r,S_{i}\right)$. Let $j\in[\ell,k]$ be the minimum integer that satisfy the following:
  1. (i)
    $j=k$,
  2. (ii)
    $G\left[P\left(r,S_{j+1}\right)\right]$ is connected,
  3. (iii)
    $G\left[P\left(r,S_{j+1}\right)\right]$ has two connected components $C_{1},C_{2}$ such that the sets $D_{1}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{1},uv\in E(G)\}$ and $D_{2}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{2},uv\in E(G)\}$ have a common vertex.
We prove the following claim.
Claim 13.1.
$d_{G}\left(x_{k},x_{j}\right)\leq k-j+4$.
Proof.
If $k=j$, then the claim is trivially true. So we assume $j<k$. Let $x^{\prime}_{k}\coloneq x_{k}$ and for each $i\in[k-1,j]$, let $x^{\prime}_{i}\in S_{i}$ denote a vertex such that $x^{\prime}_{i}x^{\prime}_{i+1}\in E(G)$. Let $R\coloneq x^{\prime}_{k}x^{\prime}_{k-1}\ldots x^{\prime}_{j}$. If $G\left[P\left(r,S_{j+1}\right)\right]$ is connected, then due to Lemma˜5(b), there is a $(x^{\prime}_{j},x_{j})$-induced path $Q^{\prime}\subseteq F$ of length at most three in $G$. Hence, $d_{G}\left(x_{k},x_{j}\right)\leq k-j+4$.
Now suppose $G\left[P\left(r,S_{j+1}\right)\right]$ has two connected components $C_{1},C_{2}$ such that the sets $D_{1}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{1},uv\in E(G)\}$ and $D_{2}\coloneq\{u\in S_{j+1}\colon\exists v\in C_{2},uv\in E(G)\}$ have a common vertex. Since $D_{1}\cap D_{2}\neq\emptyset$, due to Algorithm˜1 of Algorithm˜1, there exist $u\in S_{j+1}$ and $w,v\in S_{j}$ such that $\{x^{\prime}_{j}w,wu,uv,vx_{j}\}\subseteq E(G)$. Hence, $d_{H}\left(x^{\prime}_{j},x_{j}\right)\leq 4$ and thus, $d_{H}\left(x_{k},x_{j}\right)\leq k-j+4$.
Let $x^{\prime}_{j}=x_{j}$ and for each $i\in[\ell,j-1]$, let $x^{\prime}_{i}\in P\left(r,S_{i+1}\right)$ be a vertex such that $x^{\prime}_{i}x^{\prime}_{i+1}\in E(G)$. We prove the following claim.
Claim 13.2.
For each $i\in[\ell,j-1]$, $x_{i}x^{\prime}_{i}\in E(G)$.
Proof.
For each $i\in[\ell,j-1]$, let the two connected components of $G\left[P\left(r,S_{i+1}\right)\right]$ are $C^{1}_{i}$ and $C^{2}_{i}$. Due to Proposition˜1, both $C^{1}_{i}$ and $C^{2}_{i}$ are cliques in $G$. For each $t\in\{1,2\}$, let $D^{t}_{i}\coloneq\{u\in S_{i+1}\colon\exists v\in C^{t}_{i},uv\in E(G)\}$. By definition, $D^{1}_{i}\cap D^{2}_{i}=\emptyset$. Without loss of generality assume that $x_{i+1}\in D^{1}_{i}$. Algorithm˜1 ensures that $x_{i}\in C^{1}_{i}$. Observe that $x^{\prime}_{i+1}\in D^{1}_{i}$. Moreover, any vertex in $S_{i}$ which is adjacent to $x^{\prime}_{i+1}$ in $G$ must lie in $C^{1}_{i}$. Therefore, $x^{\prime}_{i}\in C^{1}_{i}$. Since $C^{1}_{i}$ is a clique, we have that $x_{i}x^{\prime}_{i}\in E(G)$. ∎
Let $R_{1}$ be a $(x_{k},x_{j})$-isometric path in $G$, and $R_{2}\coloneq x^{\prime}_{j}x^{\prime}_{j-1}\ldots x^{\prime}_{\ell}$. Due to ˜13.1, length of $R_{1}$ is at most kj+4k-j+4. Length of $R_{2}$ is $j$-$\ell$. Due to ˜13.2, $x_{\ell}x^{\prime}_{\ell}\in E(G)$. Hence $d_{G}\left(x,y\right)\leq d_{H}\left(x,y\right)+5$. ∎
3.3 Induced cycles in $H$
The main purpose of this section is to prove that any cluster of $G$ contains at most two vertices of an induced cycle in $H$ (see Lemma˜16). We start by proving the following.
Lemma 14.
Let $C$ be an induced cycle in $H$. For $k\geq 1$, all vertices in $V(C)\cap N_{k}(r)$ lies in the same cluster.
Proof.
Assume there exists an integer $k$ and two vertices $u_{1},u_{2}\in V(C)\cap N_{k}(r)$ such that $u_{1},u_{2}$ lie in different clusters $S_{1},S_{2}$ respectively. Let $S$ be the least common ancestor of $S_{1}$ and $S_{2}$ in $\Gamma\left({r},{G}\right)$. Observe that $S,S_{1},S_{2}$ are all distinct clusters. For $i\in\{1,2\}$, let $S^{\prime}_{i}$ denote the cluster which is a child of $S$ and lie in the $(S,S_{i})$-path in $\Gamma\left({r},{G}\right)$. Observe that $S,S^{\prime}_{1},S^{\prime}_{2}$ are distinct clusters. Since there are two $(u_{1},u_{2})$-paths in $C$ (and thus in $H$) which are internally vertex disjoint, due to ˜2, $S$ must contain two vertices $v_{1},v_{2}\in V(C)$. Observe that for $i\in\{1,2\},j\in\{1,2\}$, $v_{i}$ must have a neighbor in $C$ (and thus in $H$) that lies in $S^{\prime}_{j}$. Now applying ˜3(b) on $S^{\prime}_{1}$, we have that $G\left[P\left(r,S^{\prime}_{1}\right)\right]$ has two connected components $C_{1},D_{1}$. Similarly, applying ˜3(b) on $S^{\prime}_{2}$, we have that $G\left[P\left(r,S^{\prime}_{2}\right)\right]$ has two connected components $C_{2},D_{2}$. Without loss of generality, $v_{1}\in C_{1}\cap C_{2}$ and $v_{2}\in D_{1}\cap D_{2}$. But this contradicts Lemma˜6. ∎
Lemma 15.
Let $C$ be an induced cycle in $H$, $k=\max\{d_{H}\left(r,v\right)\colon v\in V(C)\}$ and $\ell=d_{H}\left(r,V(C)\right)$. Then, $(a)~|N_{k}(r)\cap V(C)|\leq 2$, and $(b)~|N_{\ell}(r)\cap V(C)|\leq 2$.
Proof.
Due to Lemma˜8, $k=\max\{d_{G}\left(r,v\right)\colon v\in V(C)\}$ and $\ell=d_{G}\left(r,V(C)\right)$. Let $S_{k}\subseteq N_{k}(r)$ and $S_{\ell}\subseteq N_{\ell}(r)$ denote the clusters that contain some vertices of $C$. Due to Lemma˜14, $V(C)\cap N_{k}(r)\subseteq S_{k}$ and $V(C)\cap N_{\ell}(r)\subseteq S_{\ell}$. Now, $(a)$ follows from ˜4.
To prove $(b)$, assume $\{a,b,c\}\subseteq S_{\ell}\cap V(C)$ and $F=\{uv\in E(H)\colon\{u,v\}\subseteq\{a,b,c\}\}$. By ˜5(a), $|F|\leq 1$. Due to Lemma˜14, there exists a child $S^{\prime}$ of $S_{\ell}$ in $\Gamma\left({r},{G}\right)$ such that for each $d\in\{a,b,c\}$ there exists $d^{\prime}\in S^{\prime}$ such that $dd^{\prime}\in E(H)$. This contradicts ˜3(a). ∎
Now we are ready to prove the main lemma of this section.
[…]
(a) Notations used in the proof of ˜16.1.
[…]
(b) Notations used in the proof of ˜16.2.
[…]
(c) Notations used in the proof of ˜16.3.
Figure 3: Illustration of the proof of Lemma˜16. The dashed lines indicate edges in $H$. The solid lines indicate edges in $G$. The curved lines indicate induced paths in $G$. The vertices in the shaded region indicate cliques in $G$. There are no edges in $G$ between a vertex in a rectangular region and a vertex in a elliptical region.
Lemma 16.
Let $C$ be an induced cycle in $H$ and $S$ be a cluster. Then $|S\cap V(C)|\leq 2$.
Proof.
Let $k=\max\{d_{H}\left(r,v\right)\colon v\in V(C)\}$ and $\ell=d_{H}\left(r,V(C)\right)$. Due to Lemma˜8, $k=\max\{d_{G}\left(r,v\right)\colon v\in V(C)\}$ and $\ell=d_{G}\left(r,V(C)\right)$. Due to Lemma˜14, there are clusters $S_{k}$ and $S_{\ell}$ such that $V(C)\cap N_{k}(r)\subseteq S_{k}$ and $V(C)\cap N_{k}(r)\subseteq S_{\ell}$. Due to Lemma˜15, we have that $|S_{k}\cap V(C)|\leq 2$ and $|S_{\ell}\cap V(C)|\leq 2$. Moreover, all vertices of $C$ lie in the clusters that lie in the unique $(S_{k},S_{\ell})$-path in $\Gamma\left({r},{G}\right)$.
Assume the lemma is not true, and let $k^{\prime}$ be the maximum integer such that there exists a cluster $S\subseteq N_{r}(k^{\prime})$ with $|V(C)\cap S|\geq 3$. Due to Lemma˜15, $\ell<k^{\prime}<k$. Let $T$ be the child of $S$ that lie in the unique $(S_{k},S)$-path in $\Gamma\left({r},{G}\right)$, and $U$ be the parent of $S$ in $\Gamma\left({r},{G}\right)$. Observe that $U\subseteq N_{k-1}(r)$ and $T\subseteq N_{k+1}(r)$. Due to ˜3(a), there exists at least one vertex $a\in V(C)\cap S$ such that $T$ does not contain any vertex of $C$ which is adjacent to $a$. Let $a_{1},a_{2}$ be the two neighbors of $a$ in $C$. We prove the following claim.
Claim 16.1.
$\{a_{1},a_{2}\}\cap S\neq\emptyset$.
Proof.
Assuming otherwise, let $\{a_{1},a_{2}\}\subseteq U$. See Figure˜3(a) for notations. Since $U$ is the parent of $S$ in $\Gamma\left({r},{G}\right)$, $P\left(r,S\right)\subseteq U$. Now, $P\left(r,S\right)$ contains two vertices $a_{1},a_{2}$ which is adjacent to some vertices (e.g. $a$) in $S$. Due to ˜3(b), we have that $G\left[P\left(r,S\right)\right]$ have two connected components $C_{1},C_{2}$ such that $a_{1}\in C_{1},a_{2}\in C_{2}$. Observe that, $C_{1},C_{2}$ induce cliques in $G$ and that no vertex of $U\setminus\{a_{1},a_{2}\}$ is adjacent to any vertex of $S$.
Let $x$ be a vertex in $S_{k}\cap V(C)$. Observe that there exist a $(x,a_{1})$-path $P_{1}\subseteq C$ that does not contain $a$. Similarly, there exists a $(x,a_{2})$-path $P_{2}\subseteq C$ that does not contain $a$. Observe that $P_{1}$ and $P_{2}$ only meets at $x$. Moreover, for $i\in\{1,2\}$ there must exists a vertex $b_{i}\in S\cap V(P_{i})$ such that $a_{i}b_{i}\in E(C)$. For $i\in\{1,2\}$, let $c_{i}\in V(P_{i})\setminus\{a_{i}\}$ be the neighbor of $b_{i}$ distinct from $a_{i}$. For each $i\in\{1,2\}$, observe that $c_{i}\in T$, otherwise $c_{i}$ will be adjacent to $a_{1}$ or $a_{2}$ (in $H$) contradicting the fact that $C$ is an induced cycle.
Now apply ˜1 on each of the edges in $\{aa_{1},aa_{2},b_{1}a_{1},b_{2}a_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{a_{1}^{\prime},a_{2}^{\prime},a^{\prime\prime}_{1},a^{\prime\prime}_{2}\}\subseteq U$ such that $\{aa^{\prime}_{1},aa^{\prime}_{2},b_{1}a^{\prime\prime}_{1},b_{2}a^{\prime\prime}_{2}\}\subseteq E(G)$ and $\{a^{\prime}_{1},a^{\prime\prime}_{1},a_{1}\}\subseteq C_{1}$ and $\{a^{\prime}_{2},a^{\prime\prime}_{2},a_{2}\}\subseteq C_{2}$. Similarly, apply ˜1 on each of the edges in $\{b_{1}c_{1},b_{2}c_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{b^{\prime}_{1},b^{\prime}_{2}\}\subseteq S$ such that $\{b^{\prime}_{1}c_{1},b^{\prime}_{2}c_{2},b_{1}b^{\prime}_{1},b_{2}b^{\prime}_{2}\}\subseteq E(G)$.
Since $c_{1},c_{2}$ lies in the same cluster $T\subseteq N_{k+1}(r)$, there exists a $(c_{1},c_{2})$-induced path $Q$ in $G-B_{k}\left({r}\right)$. On the other hand there exists a $(a_{1},a_{2})$-upper path (see Section˜2.2 for definition) $R$ in $G$ such that $(V(R)\setminus\{a_{1},a_{2}\})\subseteq B_{k-2}\left({r}\right)$. Now consider the vertices
$Z\coloneq V(R)\cup\{a,a_{1},a_{2},a^{\prime}_{1},a^{\prime}_{2},a^{\prime\prime}_{1},a^{\prime\prime}_{2}\}\cup\{b_{1},b_{2},b^{\prime}_{1},b^{\prime}_{2}\}\cup\{c_{1},c_{2}\}\cup V(Q)$
It can be verified that that $G\left[Z\right]$ contains a theta, long prism, pyramid, or a broken wheel as an induced subgraph, which contradicts Proposition˜2. ∎
Due to ˜16.1, we assume (without loss of generality) $a_{1}\in U,a_{2}\in S$. For $i\in\{1,2\}$, let $b_{i}$ be the neighbor of $a_{i}$ in $C$ which is distinct from $a$. Now we prove the following claim.
Claim 16.2.
The vertex $b_{2}$ does not lie in $U$.
Proof.
Assume $b_{2}\in U$. See Figure˜3(b) for notations. Due to ˜3, we have that $G\left[P\left(r,S\right)\right]$ have two connected components $A,B$ such that $a_{1}\in A,b_{2}\in B$. Moreover, $A,B$ induce cliques in $G$ and no vertex of $U\setminus\{a_{1},b_{2}\}$ is adjacent to any vertex of $S$.
Now consider a vertex $x\in V(C)\cap S_{k}$. Observe that there exist a $(x,a_{1})$-path $P_{1}\subseteq C$ that does not contain $a$ or $a_{2}$. Similarly, there exists a $(x,b_{2})$-path $P_{2}\subseteq C$ that does not contain $a,a_{1}$ or $a_{2}$. Moreovoer, $V(P_{1})\cap V(P_{2})=\{x\}$ and the edge $a_{1}b_{1}$ lies in $E(P_{1})$. Let $c_{2}$ be the neighbor of $b_{2}$ in $P_{2}$. Due to ˜3, we have that $b_{1}\in S$ and $c_{2}\in S$.
Let $d_{1}\in V(P_{1})$ be the neighbor of $b_{1}$ distinct from $a_{1}$. Let $d_{2}\in V(P_{2})$ be the neighbor of $c_{2}$ distinct from $b_{2}$. For each $i\in\{1,2\}$, observe that $d_{i}\in T$, otherwise $d_{i}$ will be adjacent to $a_{1}$ or $b_{2}$ contradicting the fact that $C$ is an induced cycle. Now apply ˜1 on each of the edges in $\{aa_{1},a_{2}b_{2},b_{1}a_{1},b_{2}c_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{a_{1}^{\prime},b_{2}^{\prime},a^{\prime\prime}_{1},b^{\prime\prime}_{2}\}\subseteq U$ such that $\{aa^{\prime}_{1},a_{2}a^{\prime}_{2},b_{1}a^{\prime\prime}_{1},c_{2}b^{\prime\prime}_{2}\}\subseteq E(G)$, $\{a^{\prime}_{1},a^{\prime\prime}_{1},a_{1}\}\subseteq A$ and $\{b^{\prime}_{2},b^{\prime\prime}_{2},b_{2}\}\subseteq B$. Similarly, apply ˜1 on edges in $\{b_{1}d_{1},c_{2}d_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{b^{\prime}_{1},c^{\prime}_{2}\}\subseteq S$ such that $\{b^{\prime}_{1}d_{1},d_{2}c^{\prime}_{2},b_{1},b^{\prime}_{1},c_{2}c^{\prime}_{2}\}\subseteq E(G)$.
Since $d_{1},d_{2}$ lies in the same cluster $T\subseteq N_{k+1}(r)$, there exists a $(d_{1},d_{2})$-induced path $Q$ in $G-B_{k}\left({r}\right)$. On the other hand there exists a $(a_{1},b_{2})$-upper path $R$ such that $(V(R)\setminus\{a_{1},a_{2}\})\subseteq B_{k-2}\left({r}\right)$. Now consider the vertices
$Z\coloneq V(R)\cup\{a,a_{1},a_{2},a^{\prime}_{1},b^{\prime}_{2},a^{\prime\prime}_{1},b^{\prime\prime}_{2}\}\cup\{b_{1},b_{2},b^{\prime}_{1},b^{\prime}_{2}\}\cup\{c_{1},c_{2}\}\cup V(Q)$
It can be verified that $G\left[Z\right]$ contains a theta, long prism, pyramid, or a broken wheel as an induced subgraph, which contradicts Proposition˜2. ∎
We also prove the following claim.
Claim 16.3.
The vertex $b_{2}$ does not lie in $T$.
Proof.
Assume $b_{2}\in T$. See Figure˜3(c) for notations. Let $z$ be the vertex in $U$ which is also adjacent to $a_{2}$ in $H$. Due to ˜5, we have that $z$ is distinct from $a_{1}$. Moreover, due to ˜3, we have that $G\left[P\left(r,S\right)\right]$ have two connected components $A,B$ such that $a_{1}\in A,z\in B$. Moreover, $A,B$ induce cliques in $G$ and no vertex of $U\setminus\{a_{1},z\}$ is adjacent to any vertex of $S$.
Now consider a vertex $x\in V(C)\cap S_{k}$. Observe that there exists a $(x,a_{1})$-path $P\subseteq C$ that does not contain $a$ or $a_{2}$. Moreover, $a_{1}b_{1}\in E(P)$ and $z\notin V(P)$. Due to ˜3, we have that $b_{1}\in S$. Let $c\in V(P)$ be the neighbor of $b_{1}$ distinct from $a_{1}$. Since $H$ is triangle-free, $c$ is not adjacent to $a_{1}$. Due to ˜5, we have $c\neq z$ and moreover, $c$ is not adjacent to $z$ in $H$. Hence, $c\notin U,c\notin S$. Therefore, $c$ must lie in $T$.
Now apply ˜1 on each of the edges in $\{aa_{1},a_{2}z,b_{1}a_{1}\}\subseteq E(H)$ to infer that there exists vertices $\{a_{1}^{\prime},z^{\prime},a^{\prime\prime}_{1}\}\subseteq U$ such that $\{aa^{\prime}_{1},a_{2}z^{\prime},b_{1}a^{\prime\prime}_{1}\}\subseteq E(G)$, $\{a^{\prime}_{1},a^{\prime\prime}_{1},a_{1}\}\subseteq A$ and $\{z,z^{\prime}\}\subseteq B$. Similarly, apply ˜1 on edges in $\{b_{1}c,a_{2}b_{2}\}\subseteq E(H)$ to infer that there exists vertices $\{b^{\prime}_{1},a^{\prime}_{2}\}\subseteq S$ such that $\{b^{\prime}_{1}c,a^{\prime}_{2}b_{2},b_{1}b^{\prime}_{1},a_{2}a^{\prime}_{2}\}\subseteq E(G)$.
Since $c,b_{2}$ lies in the same cluster $T\subseteq N_{k+1}(r)$, there exists a $(c,b_{2})$-induced path $Q$ in $G-B_{k}\left({r}\right)$. On the other hand there exists a $(a_{1},z)$-upper path $R$ such that $(V(R)\setminus\{a_{1},z\})\subseteq B_{k-2}\left({r}\right)$. Now consider the vertices
$Z\coloneq V(R)\cup\{a,a_{1},a_{2},a^{\prime}_{1},a^{\prime\prime}_{1}\}\cup\{z,z^{\prime}\}\cup\{b_{1},b_{2},b^{\prime}_{1},a^{\prime}_{2}\}\cup V(Q)$
It can be verified that $G\left[Z\right]$ contains a theta, long prism, pyramid, or a broken wheel as an induced subgraph, which contradicts Proposition˜2. ∎
Claims 16.1, 16.2, and 16.3 imply that $\{a,a_{2},b_{2}\}\subseteq S$. Since $\{aa_{2},a_{2}b_{2}\}\in E(H)$, we have a contradiction due to ˜5. ∎
3.4 Tree-width of $H$
In this section, we shall show that the output of Algorithm˜1 is a triangle-free universally signable graph and conclude using Proposition˜5. We begin by observing that $H$ is triangle-free.
Lemma 17.
The graph $H$ is triangle-free.
Proof.
Assume for contradiction that, $H$ contains a triangle $T$. Then there must exist a cluster $S$ w.r.t $r$ and an edge $uv\in E(T)$ such that $\{u,v\}\subseteq S$. This contradicts ˜5(b). ∎
Lemma 18.
The graph $H$ is theta-free.
Proof.
Assume $H$ contains a theta $T$ as an induced subgraph. Let $u,v\in V(T)$ be the vertices of degree three and let $P_{1},P_{2},P_{3}$ be the three $(u,v)$-paths of $T$. Without loss of generality, assume $d_{G}\left(r,u\right)\leq d_{G}\left(r,v\right)$ and $S$ be the cluster that contains $v$.
Suppose $u,v$ lies in different clusters. For each $i\in\{1,2,3\}$, there exist an edge $x_{i}y_{i}\in E(P_{i})$ such that $x_{i}\in S$ and $y_{i}\in P\left(r,S\right)$. If $y_{1},y_{2},y_{3}$ are distinct vertices, we have a contradiction due to ˜3(a). Therefore, we can assume without loss of generality that $y_{1}=y_{2}=u$. This implies $x_{1}$ and $x_{2}$ are distinct vertices, and that $x_{1}\neq v,x_{2}\neq v$. However, now $S$ contains three vertices from an induced cycle in $H$ (formed by $V(P_{1})\cup V(P_{2})$), which contradicts Lemma˜16.
Hence $u,v$ lies in the same cluster and let $U$ be the cluster that contains $P\left(r,S\right)$. Suppose there exists two vertices $\{v_{1},v_{2}\}\subseteq U$ such that $\{vv_{1},vv_{2}\}\subseteq E(T)$. Without loss of generality, assume that $v_{1}\in V(P_{1})$ and $v_{2}\in V(P_{2})$. Let $C$ be the induced cycle in $H$ induced by $V(P_{1})\cup V(P_{2})$. Then, for each $i\in\{1,2\}$ there exist edges $x_{i}y_{i}\in E(P_{i})$ such that $\{x_{1},x_{2}\}\subseteq U$ and $\{y_{1},y_{2}\}\subseteq S$. Applying Lemma˜16 on $C$, $S$, and $U$, we have that $\{x_{1},x_{2}\}=\{v_{1},v_{2}\}$ and $\{y_{1},y_{2}\}=\{u,v\}$. However, this contradicts ˜4. Using similar arguments, we can show that there does not exist a cluster $S^{\prime}$ in $G$ such that $P\left(r,S^{\prime}\right)\subseteq S$ and $S^{\prime}$ contains two vertices that are adjacent to $v$ in $T$. Hence, it must be the case that there exists a vertex $v_{1}\in S\setminus\{u,v\}$ such that $v,v_{1}\in V(T)$. Hence, $S$ contains three vertices $u,v,v_{1}$ from an induced cycle in $H$, which contradicts Lemma˜16. ∎
Lemma 19.
The graph $H$ is wheel-free.
Proof.
Assume that $H$ contains a wheel whose rim is $C$ and the center is $u$. Let $S$ be the cluster that contains $u$ and let $u_{1},u_{2},u_{3}$ be three neighbors of $u$ in $C$. Suppose there exists a vertex $z\in\{u_{1},u_{2},u_{3}\}$ such that $d_{G}\left(r,z\right)=d_{G}\left(r,u\right)+1$ and $T$ be the cluster that contains $z$. Note that $P\left(r,T\right)\subseteq S$. Moreover, there must be a $z^{\prime}\in\{u_{1},u_{2},u_{3}\}\setminus\{z\}$ such that $z^{\prime}\notin T$. (Otherwise, $T$ will contain three vertices of $C$, which contradicts Lemma˜16.) Suppose $d_{G}\left(r,z^{\prime}\right)=d_{G}\left(r,u\right)-1$. Then there exists two distinct $(z,z^{\prime})$-paths $Q_{1}\subseteq C,Q_{2}\subseteq C$, such that $V(Q_{1})\cap V(Q_{2})=\{z,z^{\prime}\}$. Observe that for each $i\in\{1,2\}$, there must be an edge $x_{i}y_{i}\in E(Q_{i})$ such that $x_{i}\in S$ and $y_{i}\in T$. Now, let $F\coloneq\{w\in P\left(r,T\right)\colon\exists v\in T,wv\in E(H)\}$. Observe that $\{x_{1},x_{2},u\}\subseteq F$, and hence $|F|\geq 3$, which contradicts ˜3(a). This implies $\{u_{1},u_{2},u_{3}\}\setminus\{z\}\subseteq S$, but this contradicts ˜5(a). ∎
Propositions˜5, 17, 18 and 19 implies the following.
Lemma 20.
The tree-width of $H$ is at most two.
3.5 Completion of the proof
Recall that $G$ is a $K_{2,3}$-induced minor free graph, $r$ is an arbitrary vertex of $G$ (fixed in the beginning of 3), $H$ is the graph returned by Algorithm˜1. Let $c\coloneq 2470$ and $u,v$ be any two vertices of $G$. Due to Proposition˜7, the strong isometric path complexity of $G$ is at most $71$. Due to Lemmas˜7, 12 and 10, $d_{H}\left(u,v\right)-d_{G}\left(u,v\right)\leq c$. Due to Lemmas˜17, 18 and 19 we know that $H$ is a universally signable graph, and therefore $K_{2,3}$-induced minor-free. Hence, the strong isometric path complexity of $H$ is at most $71$. Due to Lemmas˜7, 11 and 13, we have that $d_{H}\left(u,v\right)-d_{G}\left(u,v\right)\leq c$. Hence, $\left|d_{H}\left(u,v\right)-d_{G}\left(u,v\right)\right|\leq c$. Finally, Lemma˜20 asserts that $H$ has tree-width at most two. This completes the proof.
3.6 Remark on Corollary˜1
Any minimal cutset of universally signable graphs is either a clique or a stable set of size two (Proposition˜3). Therefore, when inputs of Algorithm˜1 are restricted to universally signable graphs the if condition of Algorithm˜1 can be modified as follows:
If $G\left[P\left(r,S\right)\right]$ contains at least three vertices or is an edge
This check can be done in $O(1)$-time, and the running time of Algorithm˜1 will be $O(n+m)$ where $n$ and $m$ are the number of vertices and edges of the input universally signable graph. Since the diameter of graphs of tree-width two can be found optimally in truly sub-quadratic time [7], we have the proof of Corollary˜1.
4 Conclusion
In this paper, combining the layering partition technique with the notion of strong isometric path complexity, we proved that $K_{2,3}$-induced minor-free graphs admit quasi-isometry with constant additive distortion to graphs with tree-width at most two. Minimizing the upper bound on the additive distortion is interesting. By tailoring the proof technique of Chakraborty et al. (Theorem 3, [8]) slightly and using Proposition˜1, the bound on the strong isometric path complexity of $K_{2,3}$-induced minor free can be improved. This would lead to an improvement on the bound on the additive distortion (to below $1000$). However, we believe the actual value of the optimum additive distortion consists of single digits. Providing a linear time algorithm to map $K_{2,3}$-induced minor-free graphs with additive distortion to graphs with tree-width at most two would be interesting.
Since $K_{2,r}$-induced minor-free graphs and more generally $K_{2,r}$-asymptotic minor-free graphs [22] have bounded strong isometric path complexity, it would be interesting to study if these graph classes admit quasi-isometry with additive distortion to graphs with tree-width at most $r$. It would also be interesting to study if hereditary graph classes like even-hole free graphs [29], pyramid-only graphs, etc. admit quasi-isometry with additive distortion to graphs with tree-width at most two. A result of Albrechtsen et al. [2] implies that even-hole free and pyramid-only graphs admit quasi-isometry to graphs with tree-width at most two. However, both even-hole free graphs and pyramid-only graphs have unbounded strong isometric path complexity. Therefore, our technique cannot be used directly.